A book shelf containing 20 books of which 12 are on circuit theory and 8 are on Mathematics. If three books are taken out at random, find the probability that all the three are on the same subject.
Answer by Theo(13342)

20 books.
12 on circuit theory.

8 on mathematics.

you pick 3 at random.

probability that all 3 are on circuit theory would be 12/20 * 11/19 * 10/18

probability that all 3 are on mathematics would be 8/20 * 7/19 * 6/18

probability that all 3 are either on circuit theory or on mathematics would be:

(12/20 * 11/19 * 10/18) + (8/20 * 7/19 * 6/18)

you can see this with smaller numbers.

assume 5 books total.

3 on circuit theory
2 on math

you pick 2 at random.

probability all on circuit theory is 3/5 * 2/4 = 6/20

probability all on math is 2/5 * 1/4 = 2/20

probability either all circuit theory or all math is 6/20 + 2/20 = 8/20

let abcde be the books.

possible permutations of 2 are:
ab
ac
ad
ae
ba
bc
bd
be
ca
cb
cd
ce
da
db
dc
de
ea
eb
ec
ed

assume abc are books on circuit theory

look for all permutations of ab or ac or bc

that would be ab ba ac ca bc cb

that should be 6 occurrences out of 20

assume de are books on math.

look for all permutations of de

that would be de and ed

that should be 2 occurrences out of 20

total possible occurrences is 8 / 20 which is the same as derived above.